# Generalized saturation game

### v1

@@ 3.4 Counting $P_{6}$’s
    There is a $P_{i}\in S$, where $i\in\{3,4,5\}$. Then Mini closes it to a $C_{i}$.
@@
    $S=2K_{2}$. Then Mini completes a $P_{4}$, connecting these isolated edges.
@@
    $S=K_{2}$. Mini completes a $P_{3}$, connecting one isolated vertex to $K_{2}$.
@@
    $S=\emptyset$. Mini plays an isolated $K_{2}$.
@@
At the moment when Mini has finished her move, the graph contains, some $C_{3}$s, some $C_{4}$s, some $C_{5}$s, and at most one component from the set $\{K_{2},P_{3},P_{4}\}$. Therefore, after the Max’s move the set $S$ is exactly one of the $\{\emptyset,K_{2},2K_{2},P_{3},P_{4},P_{5},P_{3}+K_{2},P_{4}+K_{2}\}$. Following her strategy, Mini ensures that there is no cycle on more than $5$ vertices in the graph. Hence, ${s(n,P_{6},S_{4})=0}$. ∎
@@ 4 ${\mathcal{C}}$-saturation games
If a tree $T$ on $n$ vertices contains at least $x$ vertices of degree at least 2, then the number of $S_{k}$’s in $T$ is at most $\binom{n-x}{k-1}$, provided $k\geq 4$.
@@
The number of $S_{k}$’s is $\sum_{i=1}^{n}\binom{d_{i}}{k-1}$, where $d_{i}$ is the degree of vertex $v_{i}\in V(T)$. For any $a\leq b$, we have $\binom{a}{k-1}+\binom{b}{k-1}\leq\binom{a-1}{k-1}+\binom{b+1}{k-1}$. Therefore, under the conditions of the lemma, $\sum_{i=1}^{n}\binom{d_{i}}{k-1}$ is maximal if $x-1$ vertices have degree 2, $n-x$ vertices have degree 1, and one vertex has degree $n-x$. This is realizable by a $P_{x+2}$ with all vertices not on the path connected to the same vertex of the path that is not end-vertex. ∎
@@
If a tree $T$ on $n$ vertices contains a double star $D_{x,y}$, then $T$ contains at least $xy+\min\{x,y\}+(n-x-y-2)-1$ copies of $P_{4}$.
@@
Let $T$ be a tree on $n\geq s\geq 4$ vertices. If $T$ contains a $P_{s}$, then the number of $P_{4}$’s that $T$ contains is at most
@@
If $d=3$, then $T=D_{x,y}$ with $x+y=n-2$ and the number of $P_{4}$’s is $xy$ which is maximized as claimed by the lemma.
If $d=4$, and so a longest path in $T$ is $u_{1}u_{2}u_{3}u_{4}u_{5}$, then we denote by $a$ the number of neigbors of $u_{3}$ and we let $N(u_{3})=\{v_{1},v_{2},\dots,v_{a}\}$. Denote by $b_{i}$ the number of neighbors of $v_{i}$ excluding $u_{3}$ for $1\leq i\leq a$. Then $1+a+\sum_{i=1}^{a}b_{i}=n$, and the number of $P_{4}$’s in $T$ is $(a-1)\sum_{i=1}^{a}b_{i}=(a-1)(n-a-1)$. This is maximized when $a=\lfloor\frac{n}{2}\rfloor$.
Finally, we consider $d\geq 5$. Let $P=u_{1}u_{2}\dots u_{s}$ be a longest path in $T$. Note that there are $s-3$ copies of $P_{4}$ on the path, and for any vertex $v\in V(T)\setminus V(P)$, there are at most 2 copies of $P_{4}$ starting at $v$ and ending on $P$. We need to show that there are at most $\lfloor\frac{n-s}{2}\rfloor\cdot\lceil\frac{n-s}{2}\rceil$ copies of $P_{4}$ with both end-vertices in $V(T)\setminus V(P)$. Let $v,v^{\prime}\in V(T)\setminus V(P)$ be two vertices not on the path that are furthest from each other.
@@
Then if $S^{r}_{T}(w)$ denotes the set of those vertices of $V(T)\setminus V(P)$ that are at distance exactly $r$ from $w$, then $S^{3}_{T}(v)\cap S^{3}_{T}(v^{\prime})=\emptyset$, unless $d_{T}(v,v^{\prime})=6$ in which case there can be at most one vertex in the intersection. We claim $\min\{|S^{3}_{T}(v)|,|S^{3}_{T}(v^{\prime})|\}\leq\lfloor\frac{n-s}{2}\rfloor$, which is trivial if the intersection is empty, while it follows from the fact that $v,v^{\prime}\notin S^{3}_{T}(v)\cup S^{3}_{T}(v^{\prime})$ if $d_{T}(v,v^{\prime})\neq 3$. We can apply induction to $T\setminus\{v^{*}\}$, where $v^{*}\in\{v,v^{\prime}\}$ with $|S^{3}_{T}(v^{*})|\leq\frac{n-s}{2}$ to obtain that the number of $P_{4}$’s in $T$ is at most $f(n-1,s)+2+\lfloor\frac{n-s}{2}\rfloor=f(n,s)$.
@@
Then consider the minimal subtree $T^{\prime}$ of $T$ containing all vertices of $V(T)\setminus V(P)$. As it has diameter 4, there exists a vertex $w$ with $d_{T^{\prime}}(w,y)\leq 2$, for every $y\in V(T^{\prime})$. Let $a$ denote the number of neighbors of $w$ in $V(T)\setminus V(P)$ and let $b$ denote the number of vertices in $V(T)\setminus V(P)$ that are at distance 2 from $w$. Then $a+b$ is either $n-s$ or $n-s-1$ depending on whether $w\in V(P)$ and the number of $P_{4}$’s with both end-vertices not on $P$ is at most $a\cdot b$ which is maximized when $|a-b|\leq 1$. ∎
@@
Now, we give a strategy for Mini. Based on Lemma 4.2, she needs to create at least $\lfloor\frac{n}{2}\rfloor$ vertices of degree at least 2 by the end of the game. If Mini starts, then she can make sure that after her second move, the game graph is a $P_{4}$. From then on, she always connects a leaf of a component to any other component thus creating a new non-leaf vertex. As there are $n-4$ more moves to be played, by the end of the game she ensures $2+\lfloor\frac{n-4}{2}\rfloor=\lfloor\frac{n}{2}\rfloor$ non-leaf vertices.
If Max starts by claiming an edge $e_{1}$, then Mini claims an edge $e_{2}$ with $e_{1}\cap e_{2}=\emptyset$. If Max claims $e_{3}$ with $e_{3}\cap(e_{1}\cup e_{2})\neq\emptyset$, then Mini creates a $P_{5}$ with her second move and from then on creates one new non-leaf vertex with each of her moves obtaining a total of $3+\lfloor\frac{n-5}{2}\rfloor\geq\lfloor\frac{n}{2}\rfloor$ non-leaves. Finally, if Max claims $e_{3}$ to obtain a matching with 3 edges, then Mini joins $e_{1},e_{2}$ to obtain a $P_{4}$. Now if Max with his next move, connects $P_{4}$ and $e_{3}$, then he creates a non-leaf, a Mini can always create a new non-leaf obtaining a total of $2+1+\lfloor\frac{n-6}{2}\rfloor\geq\lfloor\frac{n}{2}\rfloor$ non-leaves. Or Max does not connect, but then Mini with her first move can connect two leaves and make them non-leaves and by the end of the game achieves the same number of non-leaves as before. ∎
@@
First, we describe a strategy for Max. Using Lemma 4.1 we know that Max can build a balanced double star $D_{x,y}$ on $\lfloor\frac{n}{2}\rfloor+1$ vertices, regardless of whether he is the first or the second player. Therefore, $x+y=\lfloor\frac{n-2}{2}\rfloor$ and thus $x=\displaystyle{\Big{\lfloor}\frac{\lfloor\frac{n-2}{2}\rfloor}{2}\Big{\rfloor},y=\Big{\lceil}\frac{\lfloor\frac{n-2}{2}\rfloor}{2}\Big{\rceil}}$. Lemma 4.3 yields that the number of $P_{4}$’s is at least $xy+\min\{x,y\}+(n-x-y-2)-1$. Plugging in we get that ${s(n,\#P_{4},{\mathcal{C}})\geq\frac{n^{2}}{16}+\frac{n}{8}-\frac{27}{16}}$.
@@ 5 $P_{5}$-saturation games
    $C$ is a $K_{2}$. Max tries to claim an isolated $K_{2}$. If that is not possible, he creates a $P_{3}$ connecting $C$ with one isolated vertex and then proceeds to Stage 2.
@@
    $C$ is a $P_{3}$. Max claims the edge that completes the triangle.
@@
      If $C$ is a $P_{4}$, Max claims the edge that completes the $C_{4}$.
@@
      Else, if there is one isolated vertex, and on top of that either an isolated edge or an isolated triangle, Max claims the edge that creates, respectively, a $P_{3}$ or a triangle with one appended edge, which in this proof will denote by $G_{4}^{1}$. Then he proceeds to Stage 2.
@@
Later in Stage 2, the only components that can change the number of vertices are pairs of isolated $K_{2}$’s that can be transformed into a $K_{4}$ by the end of the game. Therefore, the worst scenario for Max is that at the end of the game, there is one component on 5 vertices, one isolated edge that cannot be added to any of the components, and all of the remaining components are triangles. This implies
$s_{2}(n,K_{3},P_{5})\geq\frac{n-4}{3}.$
@@
    Then Mini claims an edge that connects it with an isolated vertex, creating a $P_{3}$.
@@
During Stage 1, after Mini has finished her move, the graph consists of the following connected components: at most one $K_{2}$, some $P_{3}$s, some $K_{3}$s, and some connected components on at least five vertices.
@@
If there is one isolated vertex that means that each of the connected components is on three vertices or at least five vertices. The worst case for Mini is when each of the connected components is on three vertices, then it is possible that at most one of them becomes $K_{4}$, by the end of the game. That implies $s_{2}(n,K_{3},P_{5})\leq\frac{n-4}{3}+4$.
@@
First, we obtain a lower bound by describing a strategy of Max. A component is good if it contains a $C_{4}$. At the end of the game, a good component must be a $K_{4}$. Note that a $K_{4}$ component contains $12$ $P_{4}$’s.
@@
  • if Mini extends an edge to a $P_{3}$, Max extends it to a $P_{4}$,
@@
  • if all non-isolated vertices and non-good components are $P_{4}$’s and edges, then Max prefers to create a $C_{4}$, if these components form a matching, then Max claims an isolated edge.
@@
Next, we consider Mini’s strategy to obtain an upper bound on $s(n,\#P_{4},P_{5})$. A component containing a $C_{4}$ is still good. Mini makes sure that after her move, no $P_{4}$-component and no $P_{3}$-component exists as follows:
  • if Max extends an edge to a $P_{3}$, she closes it to a triangle,
  • if Max connects two edges to a $P_{4}$ or if Max connects an isolated vertex to a triangle, she closes the component a good one,
@@
Mini cannot follow this strategy once there remains at most one isolated vertex. In that case, she connects this isolated vertex to an isolated edge or a triangle. If a $P_{3}$ is created, then this path can be connected to an edge yielding a $D_{1,2}$ containing 2 copies of $P_{4}$s. All other components will either be $K_{3}$’s or $K_{4}$’s at the end of the game, so $s(n,\#P_{4},P_{5})\leq 3n$. ∎

### v2

@@ 3.4 Counting $P_{6}$’s
    $S$ contains a copy of $P_{i}$, where $i\in\{3,4,5\}$. Then Mini closes it to a copy of $C_{i}$.
@@
    $S$ consists of two independent edges. Then Mini completes a copy of $P_{4}$, connecting these isolated edges.
@@
    $S$ consists of a single edge. Mini completes a copy of $P_{3}$, connecting one isolated vertex to $K_{2}$.
@@
    $S$ is an empty set. Mini plays an isolated $K_{2}$.
@@
At the moment when Mini has finished her move, the graph contains, some copies of $C_{3}$, some copies of $C_{4}$, some copies of $C_{5}$, and at most one component from the set $\{K_{2},P_{3},P_{4}\}$. Therefore, after Max’s move the set $S$ is exactly one of the $\{\emptyset,K_{2},2K_{2},P_{3},P_{4},P_{5},P_{3}+K_{2},P_{4}+K_{2}\}$. Following her strategy, Mini ensures that there is no cycle on more than $5$ vertices in the graph. Hence, ${s(n,\#P_{6},S_{4})=0}$. ∎
@@ 4 ${\mathcal{C}}$-saturation games
If a tree $T$ on $n$ vertices contains at least $x$ vertices of degree at least 2, then the number of copies of $S_{k}$ in $T$ is at most $\binom{n-x}{k-1}$, provided $k\geq 4$.
@@
The number of copies of $S_{k}$ is $\sum_{i=1}^{n}\binom{d_{i}}{k-1}$, where $d_{i}$ is the degree of vertex $v_{i}\in V(T)$. For any $a\leq b$, we have $\binom{a}{k-1}+\binom{b}{k-1}\leq\binom{a-1}{k-1}+\binom{b+1}{k-1}$. Therefore, under the conditions of the lemma, $\sum_{i=1}^{n}\binom{d_{i}}{k-1}$ is maximal if $x-1$ vertices have degree 2, $n-x$ vertices have degree 1, and one vertex has degree $n-x$. This is realizable by a $P_{x+2}$ with all vertices not on the path connected to the same vertex of the path that is not end-vertex. ∎
@@
If a tree $T\neq D_{x,y}$ on $n$ vertices contains a double star $D_{x,y}$, then $T$ contains at least $xy+\min\{x,y\}+(n-x-y-2)-1$ copies of $P_{4}$.
@@
Let $T$ be a tree on $n\geq s\geq 4$ vertices. If $T$ contains a copy of $P_{s}$, then the number of copies of $P_{4}$ that $T$ contains is at most
@@
If $d=3$, then $T=D_{x,y}$ with $x+y=n-2$ and the number of copies of $P_{4}$ is $xy$ which is maximized as claimed by the lemma.
If $d=4$, and so a longest path in $T$ is $u_{1}u_{2}u_{3}u_{4}u_{5}$, then we denote by $a$ the number of neigbors of $u_{3}$ and we let $N(u_{3})=\{v_{1},v_{2},\dots,v_{a}\}$. Denote by $b_{i}$ the number of neighbors of $v_{i}$ excluding $u_{3}$ for $1\leq i\leq a$. Then $1+a+\sum_{i=1}^{a}b_{i}=n$, and the number of copies of $P_{4}$ in $T$ is $(a-1)\sum_{i=1}^{a}b_{i}=(a-1)(n-a-1)$. This is maximized when $a=\lfloor\frac{n}{2}\rfloor$.
Finally, we consider $d\geq 5$. Let $P=u_{1}u_{2}\dots u_{s}$ be a longest path in $T$. Note that there are $s-3$ copies of $P_{4}$ on the path, and for any vertex $v\in V(T)\setminus V(P)$, there are at most 2 copies of $P_{4}$ starting at $v$ and ending on $P$. We need to show that there are at most $\lfloor\frac{n-s}{2}\rfloor\cdot\lceil\frac{n-s}{2}\rceil$ copies of $P_{4}$ with both end-vertices in $V(T)\setminus V(P)$. Let $v,v^{\prime}\in V(T)\setminus V(P)$ be two vertices not on the path that are furthest from each other. Suppose first that $v$ is not a leaf in $T$. Then $v$ has to be the only neighbor of a vertex of $P$ in $V(T)\setminus V(P)$. But then all copies of $P_{4}$ with both end-vertices in $V(T)\setminus V(P)$ lie completely in $V(T)\setminus V(P)$. So, the number of such $P_{4}$’s, by induction, is at most $f(n-s,s^{\prime})\leq f(n-s,4)=\lfloor\frac{n-s-2}{2}\rfloor\cdot\lceil\frac{n-s-2}{2}\rceil$ as claimed. Now we can assume that both $v,v^{\prime}$ are leaves of $T$.
@@
Then if $S^{r}_{T}(w)$ denotes the set of those vertices of $V(T)\setminus V(P)$ that are at distance exactly $r$ from $w$, then $S^{3}_{T}(v)\cap S^{3}_{T}(v^{\prime})=\emptyset$, unless $d_{T}(v,v^{\prime})=6$ in which case there can be at most one vertex in the intersection (or if $D_{T}(v,v^{\prime})=4$, but that is Case B). We claim $\min\{|S^{3}_{T}(v)|,|S^{3}_{T}(v^{\prime})|\}\leq\lfloor\frac{n-s}{2}\rfloor$, which is trivial if the intersection is empty, while it follows from the fact that $v,v^{\prime}\notin S^{3}_{T}(v)\cup S^{3}_{T}(v^{\prime})$ if $d_{T}(v,v^{\prime})\neq 3$. We can apply induction to $T\setminus\{v^{*}\}$, where $v^{*}\in\{v,v^{\prime}\}$ with $|S^{3}_{T}(v^{*})|\leq\frac{n-s}{2}$ to obtain that the number of copies of $P_{4}$ in $T$ is at most $f(n-1,s)+2+\lfloor\frac{n-s}{2}\rfloor=f(n,s)$.
@@
Then consider the minimal subtree $T^{\prime}$ of $T$ containing all vertices of $V(T)\setminus V(P)$. As it has diameter 4, there exists a vertex $w$ with $d_{T^{\prime}}(w,y)\leq 2$, for every $y\in V(T^{\prime})$. Let $a$ denote the number of neighbors of $w$ in $V(T)\setminus V(P)$ and let $b$ denote the number of vertices in $V(T)\setminus V(P)$ that are at distance 2 from $w$. Then $a+b$ is either $n-s$ or $n-s-1$ depending on whether $w\in V(P)$ and the number of copies of $P_{4}$ with both end-vertices not on $P$ is at most $a\cdot b$ which is maximized when $|a-b|\leq 1$. ∎
@@
Now, we give a strategy for Mini. Based on Lemma 4.2, she needs to create at least $\lfloor\frac{n}{2}\rfloor$ vertices of degree at least 2 by the end of the game. If Mini starts, then she can make sure that after her second move, the game graph is a copy of $P_{4}$. From then on, she always connects a leaf of a component to any other component thus creating a new non-leaf vertex. As there are $n-4$ more moves to be played, by the end of the game she ensures $2+\lfloor\frac{n-4}{2}\rfloor=\lfloor\frac{n}{2}\rfloor$ non-leaf vertices.
If Max starts by claiming an edge $e_{1}$, then Mini claims an edge $e_{2}$ with $e_{1}\cap e_{2}=\emptyset$. If Max claims $e_{3}$ with $e_{3}\cap(e_{1}\cup e_{2})\neq\emptyset$, then Mini creates a copy of $P_{5}$ with her second move and from then on creates one new non-leaf vertex with each of her moves obtaining a total of $3+\lfloor\frac{n-5}{2}\rfloor\geq\lfloor\frac{n}{2}\rfloor$ non-leaves. Finally, if Max claims $e_{3}$ to obtain a matching with 3 edges, then Mini joins $e_{1},e_{2}$ to obtain a copy of $P_{4}$. Now if Max, with his next move, connects a copy of $P_{4}$ and $e_{3}$, then he creates a non-leaf, and Mini can always create a new non-leaf obtaining a total of $2+1+\lfloor\frac{n-6}{2}\rfloor\geq\lfloor\frac{n}{2}\rfloor$ non-leaves. Or Max does not connect, but then Mini with her first move can connect two leaves and make them non-leaves and by the end of the game achieves the same number of non-leaves as before. ∎
@@
First, we describe a strategy for Max. Using Lemma 4.1 we know that Max can build a balanced double star $D_{x,y}$ on $\lfloor\frac{n}{2}\rfloor+1$ vertices, regardless of whether he is the first or the second player. Therefore, $x+y=\lfloor\frac{n-2}{2}\rfloor$ and thus $x=\displaystyle{\Big{\lfloor}\frac{\lfloor\frac{n-2}{2}\rfloor}{2}\Big{\rfloor},y=\Big{\lceil}\frac{\lfloor\frac{n-2}{2}\rfloor}{2}\Big{\rceil}}$. Lemma 4.3 yields that the number of copies of $P_{4}$ is at least $xy+\min\{x,y\}+(n-x-y-2)-1$. Plugging in we get that ${s(n,\#P_{4},{\mathcal{C}})\geq\frac{n^{2}}{16}+\frac{n}{8}-\frac{27}{16}}$.
@@ 5 $P_{5}$-saturation games
    $C$ is a $K_{2}$. Max tries to claim an isolated $K_{2}$. If that is not possible, he creates a copy of $P_{3}$ connecting $C$ with one isolated vertex and then proceeds to Stage 2.
@@
    $C$ is a copy of $P_{3}$. Max claims the edge that completes the triangle.
@@
      If $C$ is a copy of $P_{4}$, Max claims the edge that completes the copy of $C_{4}$.
@@
      Else, if there is one isolated vertex, and on top of that either an isolated edge or an isolated triangle, Max claims the edge that creates, respectively, a copy of $P_{3}$ or a triangle with one appended edge, which in this proof we will denote by $G_{4}^{1}$. Then he proceeds to Stage 2.
@@
Later in Stage 2, the only components that can change the number of vertices are pairs of isolated $K_{2}$’s that can be transformed into a copy of $K_{4}$ by the end of the game. Therefore, the worst scenario for Max is that at the end of the game, there is one component on 5 vertices, one isolated edge that cannot be added to any of the components, and all of the remaining components are triangles. This implies
$s_{2}(n,\#K_{3},P_{5})\geq\frac{n-4}{3}.$
@@
    Then Mini claims an edge that connects it with an isolated vertex, creating a copy of $P_{3}$.
@@
During Stage 1, after Mini has finished her move, the graph consists of the following connected components: at most one $K_{2}$, some $P_{3}$’s, some $K_{3}$’s, and some connected components on at least five vertices.
@@
If there is one isolated vertex that means that each of the connected components is on three vertices or at least five vertices. The worst case for Mini is when each of the connected components is on three vertices, then it is possible that at most one of them becomes $K_{4}$, by the end of the game. That implies $s_{2}(n,\#K_{3},P_{5})\leq\frac{n-4}{3}+4$.
@@
First, we obtain a lower bound by describing a strategy of Max. A component is good if it contains a copy of $C_{4}$. At the end of the game, a good component must be a $K_{4}$. Note that a $K_{4}$ component contains $12$ copies of $P_{4}$.
@@
  • if Mini extends an edge to a copy of $P_{3}$, Max extends it to a copy of $P_{4}$,
@@
  • if all non-isolated vertices and non-good components are $P_{4}$’s and edges, then Max prefers to create a copy of $C_{4}$, if these components form a matching, then Max claims an isolated edge.
@@
Next, we consider Mini’s strategy to obtain an upper bound on $s(n,\#P_{4},P_{5})$. A component containing a copy of $C_{4}$ is still good. Mini makes sure that after her move, no $P_{4}$-component and no $P_{3}$-component exists as follows:
  • if Max extends an edge to a copy of $P_{3}$, she closes it to a triangle,
  • if Max connects two edges to a copy of $P_{4}$ or if Max connects an isolated vertex to a triangle, she closes the component a good one,
@@
Mini cannot follow this strategy once there remains at most one isolated vertex. In that case, she connects this isolated vertex to an isolated edge or a triangle. If a copy of $P_{3}$ is created, then this path can be connected to an edge yielding a $D_{1,2}$ containing 2 copies of $P_{4}$. All other components will either be $K_{3}$’s or $K_{4}$’s at the end of the game, so $s(n,\#P_{4},P_{5})\leq 3n$. ∎